5:00 pm - Sunday December 10, 2017

Waldo Fire: 70% Contained

Waldo Canyon Fire Update

Fire Facts:

Date started: June 23, 2012 Number of Personnel: 1,523
Location: West of Colorado Springs Crews:  38
Size:  17,920 acres Engines:  82
Percent Contained: 70%

Estimated Containment: 7/12/12

Dozers:  15

Water Tenders:  15

Cause: Under investigation Helicopters:   3 Type 12 Type 2 and  3 Type 3
Structures threatened:  3,357 residences and 50 commercial structures


Injuries to Date: 3

Cost to Date: $12,433,656

Structures lost  347





Carol Lakes and the Farish Memorial Recreation Area are open to the public.  Businesses that have approved to open are the Cave of the Winds and Cliff Dwellings.  The fire area remains closed to the public.


Do you have video clips or cellphone footage of the Waldo Canyon Fire?  Incident Commander, Rich Harvey, asks anyone with footage of the fire to contact the team. Our videographer is especially looking for footage from June 26, when the fire grew rapidly. Drop a CD or DVD off with Security at the Incident Command Post entrance at Holmes Middle School, 2455 Mesa Rd., Colorado Springs, CO 80904. Or e-mail a link or clip to cineshine@aol.com. Or call 719-328-4334.


Today’s Update:


Today, thunderstorms produced gusty winds over the fire. An island of fuel within the fire perimeter produced the most active burning. Smoke continued to be visible on the west side of the fire. The public is reminded that smoke is also expected to be visible for several days from the western perimeter, as the fire is still smoldering. Construction continued on a contingency line north of the fire along Mt. Herman Road.


With a decrease in fire activity and significant gain in containment today, release of firefighting resources has begun.  There will be a dramatic reduction in resources tomorrow as they are released to assist other higher priority fires across the nation.


Tomorrow is expected to be slightly warmer and dryer, with isolated thunderstorms developing later in the day. Firefighters’ primary focus continues to be the northwest perimeter of the fire where heat remains a burnout conducted two days ago. Other crews continue to clear out snags along Rampart Range Rd. and in the Eagle Lake area.


To the north, crews will secure and patrol spots north of Rampart Range Rd. and improve direct handline south of West Monument Creek.  On the eastern flank, crews will work to ensure the Blodgett Peak and West Monument Creek drainages are secure and will continue rehabilitation of control lines.  Mop up will continue on all flanks of the fire where feasible.



Evacuations and Closures


Evacuations are still in effect.


For more information, go to http://springsgov.com or http://www.elpasoco.com.





Filed in: community / comunidad